in an elevator moving up with an acceleration g the force exerted on the floor by a passenger of mass M is a.Mg b.1/2Mg c.zero d.2Mg

Grade:12th pass

2 Answers

Arun
6 years ago
If the Elevator is moving up with acceleration g, same will be the acceleration for the Person inside (as we see from the ground).
The person's weight is mg downwards, the only force that is acting on the Person is the force from the lift floor. Let that reaction force be R. This R has to overcome his weight and still accelerate him to g,
We have R - mg = ma (and here 'a' is given as 'g')
therefore, R = 2mg.
Khimraj
6 years ago
Draw the free body diagram of paasenger. there will be one downward foce = M*geff.
Here geff = g+g = 2g.
So force exerted on the floor by a passenger will be 2Mg.